What weight of strontium nitrate would you need to prepare 250 mLs of a 0.726 M solution?
1
Expert's answer
2020-05-16T13:40:47-0400

1.Find the amount of substance

"c=\frac{n}{V}"

"0.726=\frac{n}{0,25}"

n=0.1815

2.Find the mass

"m=M\times n=0.1815*211,628=38.41"